`what does c stand for in the inequality |5c+4| is greater than or equal to 7''.

Answer by stanbon(75887)

what does c stand for in the inequality |5c+4| is greater than or equal to 7''.
----
5c+4 >= 7 or 5c+4 <= -7
5c >= 3 or 5c <= -11
c >= 3/5 or c <= -11/5
==========================
Cheers,
Stan H.
=============